OpenStudy (anonymous):

In a collection of dimes and pennies, there are 37 coins. The total value of the coins is $2.53. How many coins of each type are there?

OpenStudy (anonymous):

here let there be x dimes and y pennies then we have x+y=37 and 10x+y=253 subtracting the two equation we have 9x=253-37=216 hence x=216/9=24 thus there are 24 dimes and 37-24=13 pennies